# Hier-SLAM++: Neuro-Symbolic Semantic SLAM with a Hierarchically Categorical Gaussian Splatting

### v1

@@ III-A1 Tree Parametrization
which corresponds to the root-to-leaf path: $g^{i}=v_{0}^{i}\xrightarrow{e_{0}}v_{1}^{i}\xrightarrow{e_{1}}\cdots\xrightarrow{e_{L-2}}v_{L-1}^{i}\xrightarrow{e_{L-1}}v_{L}^{i}$, as illustrated in Fig. 3. Take a semantic class $g=\texttt{Bed}$ as an example, a 5-level tree symbolic coding can be represented as: $\left\{v_{0}:\texttt{LargeItems}\right\}\rightarrow\left\{v_{1}:\texttt{Furnishings}\right\}\rightarrow\left\{v_{2}:\texttt{BedroomFurniture}\right\}\rightarrow\left\{v_{3}:\texttt{Rectangular}\right\}\rightarrow\left\{v_{4}:\texttt{Bed}\right\}$. The hierarchical attributes and properties of class $g^{i}$ are represented by the symbolic node information $v_{l}^{i}$, while its relationships such as ‘include’ and ‘possessing’ are represented by the edge information $e_{m}^{i}:\rightarrow$. In this way, every semantic class can be coded in a progressive, symbolic hierarchical manner, incorporating both semantic and geometric perspectives. Moreover, the standard flat representation can be seen as a single-level tree coding from the hierarchical viewpoint. Herein, we shall omit the superscripts (class indexes) of variables for clarity and compact notation.
III-A2 Tree Generation with LLMs and 3D Generative Models
@@
To represent semantic information both accurately and comprehensively, an effective tree structure should incorporate information from both semantic and geometric perspectives, as these are the two crucial types of information in an unknown environment, which are essential for achieving a comprehensive understanding of the global 3D environment. For capturing semantic meanings, such as functional relationships and summary description labels, we use Large Language Models (LLMs), specifically GPT-4 Turbo [39], known for its strong commonsense reasoning and language understanding capabilities. On the other hand, to capture general geometric information without dataset limitations, we employ 3D Generative Models, specifically text-to-3D models, as they have the capability to generate generalized and unified 3D objects with accurate shape information based on text prompts after training. We adopt the text-to-3D approach MeshGPT [26], which learns 3D symbolic embeddings with the local mesh geometry and topology of 3D assets, making it well-suited for our purpose. However, the text-to-3D method only generates 3D objects within a unified coordinate system, omitting size information. To address this issue, we complement the geometric shape information with the size attributes derived from LLMs, to capture the comprehensive geometric information for each semantic class. The leverage of both LLMs and 3D Generative Models enables the creation of an elaborate and well-structured hierarchical tree representation, where each level reflects a distinct aspect of a class’s attributes, enhancing the depth and utility of the global 3D understanding. The following part provides a detailed explanation of the tree generation process.
@@
As illustrated in Fig. 3, the tree generation process follows a top-to-bottom manner. At the upper level, the initial grouping divides all classes into several subgroups based on their typical physical sizes in scenes, obtained through prompt queries to the LLM:
@@
where $\mathcal{L}^{\text{size}}$ represents the LLMs usage for the size message, and the subscript in $\{v,e\}_{0}$ represents the generation at the initial levels. The clustering results categorize classes into groups such as small items, medium items, and large items. We provide details of the specific prompt used for the LLM in Appendix 1.
Next, the classes within each group are used as input prompts to LLMs for further clustering based on their semantic relationships, particularly their functional properties:
@@
where $\mathcal{L}^{\text{func}}$ denotes the LLMs clustering based on functional relationships, and the subscript of $\{v,e\}_{1}$ stands for the following levels generation. We demonstrate the detail prompt used for the LLM in Appendix 1.
After multiple levels generation using LLMs, 3D Generative Models are employed to cluster classes based on their geometric shapes. Specifically, we adopt MeshGPT [26] to capture the geometric information for each semantic class. MeshGPT [26] is a text-to-3D model that encodes object shape and internal topology relationships into latent quantized embeddings, forming a symbolic 3D representation in the shape space. Given a text prompt (i.e., the class name) as input to the pre-trained MeshGPT, we extract the latent quantized embeddings to represent the shape information for the class:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{z}=\mathcal{E}\{g\}\in\mathbb{N}^{F\times 2}}}$ (4)
where $g$ is the class name provided as input. $\mathcal{E}$ represents the pre-trained MeshGPT encoder, and $\bm{z}$ is the quantized geometric embedding extracted from the pre-trained encoder. For each object’s embedding, $F$ is the number of triangle faces used in the mesh representation. Since the number of triangle faces varies across different semantic objects, i.e., different values of $F$, we first compute the average shape embedding $\bm{z}_{\text{avg}}\in\mathbb{R}^{1\times 2}$ for each object. The averaged embeddings from all classes are then concatenated and subsequently normalized to a standard distribution (zero mean and unit variance) to ensure consistency across classes:
@@
where $\mathcal{N}$ and $\mathcal{C}$ denote the normalization and concatenation operations, respectively. Next, the normalized embeddings $\tilde{\bm{z}}$ are clustered using K-means++, denoted by the operator $\mathcal{K}$. To address the lack of descriptive labels for each clustering result, we further employ LLMs to summarize and refine the clustering outcomes:
@@
where $\mathcal{L}^{\text{sum}}$ represents the use of LLMs for summarization. The generated descriptive labels, such as ”box”, ”flat”, ”soft”, summarize the common geometric attributes within each grouping result, further enhancing the interpretability of the hierarchical tree. The details are illustrated in Appendix 1.
Whenever LLMs are employed, we ensure balanced, meaningful, and descriptive outputs through carefully designed prompt inputs, as detailed in Appendix 1. Furthermore, an additional LLM validator [25] is utilized after each generation level to verify clustering completeness and accuracy, ensuring that no classes are missing or incorrectly included (see Appendix 1 for detail). Finally, a global tree evaluation is conducted to detect and resolve any duplicates or omissions across tree levels, ensuring a fully balanced and complete hierarchical tree structure. We use manual inspection as the final verification to ensure the correctness and rationality of the holistically generated tree.
@@
Figure 3: Visualization of the hierarchical tree structure and semantic embedding. For tree representation, each semantic class is expressed hierarchically as $g=\{v_{l},e_{m}\}$, where edges $e_{m}$ are depicted as lines connecting tree nodes. The nodes within the same level are illustrated via the same color. For tree generation, we utilize both LLMs and 3D Generative Models to extract and group messages. For hierarchical semantic embedding, we represent semantic information using a root-to-leaf symbolic path. We propose two types of representations: the one-hot representation and the binary representation, which can reduce the original dimensionality by up to $O(\log^{2}N)$.
III-B Hierarchical optimization
@@
Based on the generated hierarchical tree structure, the semantic information is represented as compact hierarchical symbolic nodes, and further integrated into the 3D Gaussian primitives, learned in an end-to-end manner with the input frame stream. To ensure accurate semantic understanding, we introduce a semantic hierarchical loss, which consists of Inter-level and Cross-level losses. These losses are incorporated into the optimization process during the online SLAM operation. The following sections provide a detailed explanation of the proposed tree encoding and the loss calculation.
III-B1 Tree Encoding
As introduced in Eq. (1), the semantic class in Hier-SLAM++ is represented hierarchically in a root-to-leaf symbolic manner based on the generated tree. This effectively compacts semantic information, facilitates SLAM joint optimization, and provides valuable scalability for semantic understanding in unknown environments. Specifically, we propose two types of hierarchical semantic embedding approaches, both of which compact the original semantic information.
@@ III-B1 Tree Encoding
One-hot representation: As demonstrated in Fig. 3, the hierarchical semantic embedding $\bm{h}$ is composed of the one-hot embeddings across all levels:
@@
where $l$ represents the $l$-th level of the tree, and $\mathcal{C}$ denotes the concatenation operation. For each level’s representation $\bm{h}_{l}$, we use $n$-dimensional one-hot codes (in colored boxes) to encode the symbolic nodes (in colored circles) from the tree of this level, as shown in Fig. 3. The dimension $n$ corresponds to the maximum number of nodes at the $l$-th level. Using the one-hot representation, the overall dimension of the semantic embedding is the sum of the dimensions across all levels, given by $N=\sum n$. This enable a maximum reduction of up to $O(\log N)$ compared to the original dimension $N_{\text{original}}$ in the flat representation. This hierarchical representation effectively compacts the semantic embedding, reducing storage requirements and facilitating joint optimization.
Binary representation: We propose a further compact version of our hierarchical semantic representation in this section. The binary representation of the hierarchical semantic embedding $\bm{b}$ consists of the binary embeddings $\bm{b}_{l}$ across all tree levels:
$\bm{b}=\mathcal{C}(\bm{b}_{l})\in\mathbb{B}^{K},\quad\bm{b}_{l}\in\mathbb{B}^{k}$ (8)
where $\bm{b}_{l}$ is the binary representation converted from the one-hot semantic embedding $\bm{h}_{l}$ at each level: $\bm{h}_{l}\rightarrow\bm{b}_{l}$, as depicted in the last row example in Fig. 3. For instance, if the one-hot embedding has a dimension of $n=8$, this corresponds to a maximum of $8$ unique codes at this level (covering values from 00 to $7$). After binarization, the dimension of the compact version is reduced to $k=3$, since $82^{3}=$ 8. By leveraging binary representation, the hierarchical coding dimension at each level is reduced to $k=\log n$, compared to the one-hot encoding with dimension $n$. Consequently, the total semantic encoding dimension across all levels is given by $K=\sum k$, resulting in a more compact representation.
@@ III-B2 Loss Calculation
To fully optimize the hierarchical semantic message both within each level and across different levels of the tree, the proposed hierarchical loss consists of Inter-level loss $L_{\text{Inter}}$ and Cross-level loss $L_{\text{Cross}}$:
$L_{\text{{Semantic}}}=\omega_{1}L_{\text{{Inter}}}+\omega_{2}L_{\text{{Cross}}}$ (9)
where $\omega_{1}$ and $\omega_{2}$ are balancing coefficients for each loss term. Specifically, the Inter-level loss $L_{\text{Inter}}$ is employed within each level’s embedding $\bm{h}^{l}$ to ensure the optimization of semantic information within each level:
$L_{\text{{Inter}}}=\sum_{l=0}^{L}L_{\text{{ce}}}(\mathcal{S}(\bm{h}^{l}),\mathcal{P}^{l})$ (10)
where $L_{\text{{ce}}}$ indicates the cross-entropy semantic loss, and $\mathcal{P}^{l}$ denotes the semantic ground truth at level $l$. $\mathcal{S}$ stands for the Softmax operation, converting embeddings into probabilities. For our proposed compact version with binary representation $\bm{b}^{l}$, we replace the original cross-entropy loss with binary cross-entropy loss $L_{\text{bce}}$ for the binary Inter-level loss calculation:
@@
To enforce accurate understanding across tree levels, we introduce the Cross-level loss, which is computed using a semantic decoder based on the holistic hierarchical codes $\bm{h}$. Specifically, we first perform a 2D convolution operation on the semantic embedding, following an activation layer (i.e., ReLU). Then, a second 2D convolution layer is applied to map the hidden embedding into the flat coding, followed by a softmax operation to convert the embeddings into probabilities. The overall Cross-level loss $L_{\text{Cross}}$ is defined as:
$L_{\text{Cross}}=L_{\text{ce}}\big{(}\mathcal{S}(\text{decoder}(\bm{h})),\mathcal{P}\big{)},$ (12)
where $\mathcal{P}$ represents the semantic ground truth, and the definition of $\text{decoder}(\bm{h})$ is:

### v2

@@ III-A1 Hierarchical Tree Parametrization
which corresponds to a hierarchical path for the semantic meaning: $g^{i}=v_{0}^{i}\xrightarrow{e_{0}}v_{1}^{i}\xrightarrow{e_{1}}\cdots\xrightarrow{e_{L-2}}v_{L-1}^{i}\xrightarrow{e_{L-1}}v_{L}^{i}$. For example, the 5-level hierarchical symbolic representation of the class Bed can be expressed as: $g^{\texttt{bed}}=\left\{v_{0}:\texttt{LargeItems}\right\}\rightarrow\left\{v_{1}:\texttt{Furnishings}\right\}\rightarrow\left\{v_{2}:\texttt{BedroomFurniture}\right\}\rightarrow\left\{v_{3}:\texttt{Rectangular}\right\}\rightarrow\left\{v_{4}:\texttt{Bed}\right\}$. In this representation, the original semantic label $g^{i}$ is abstracted as a hierarchical path from a coarse category to a fine-grained concept. At each level, symbolic nodes $v_{l}^{i}$ represent the semantic attributes of the class, while inter-level relationships, such as inclusion and possession, are described by the edge elements $e_{m}^{i}$. As will be demonstrated in Section III-B, with a carefully structured design, the proposed hierarchical representation yields a compact and computationally efficient representation. It allows each semantic class to be expressed in a progressive, interpretable, and symbolic manner, incorporating both semantic abstraction and geometric structure. Moreover, from a hierarchical perspective, the original flat representation can be regarded as a degenerate case of a single-level tree, enabling a unified formulation. For clarity and notational simplicity, we omit the superscripts (i.e., class indices) of variables in the remainder of this section.
III-A2 Tree Generation via LLMs and 3D Generative Models
@@
To represent semantic information in a reasonable, comprehensive, and compact manner, the tree structure should integrate both semantic and geometric cues, two fundamental sources of information in unknown environments that are essential for achieving a holistic understanding of the global 3D scene. To extract high-level semantic concepts, including functional relationships and abstract descriptors, we employ large language models (LLMs), specifically GPT-4 Turbo [69], which is well known for its strong commonsense reasoning and language understanding abilities. On the other hand, to capture general geometric information beyond dataset-specific constraints, we employ 3D generative models, particularly text-to-3D frameworks. These models are capable of producing diverse and geometrically accurate 3D objects from textual prompts after appropriate training, offering a unified and generalizable geometric representation. Specifically, we adopt MeshGPT [23], a text-to-3D model that learns symbolic 3D embeddings for the local mesh geometry and topology of each object. This property makes it particularly well-suited for our goal of constructing geometry-aware hierarchical semantic structures. However, the text-to-3D method generates 3D objects within a unified coordinate system but omits explicit size information. To address this limitation, we complement the generated geometric shapes with size attributes inferred from large language models (LLMs), enabling a more comprehensive geometric representation for each semantic class. Leveraging both large language models (LLMs) and 3D generative models enables the construction of an elaborate and well-structured hierarchical tree, where each level captures a distinct aspect of a class’s attributes—semantic, functional, or geometric, which further enriches the depth and expressiveness of global 3D scene understanding. The remainder of this section presents a detailed description of the hierarchical tree generation process.
@@
The overall process of hierarchical tree generation is illustrated in Fig. 3. At the top level, the initial grouping divides all semantic classes into several subgroups based on their typical physical sizes in real-world scenes, which are obtained through prompt-based queries to the LLM. The generation of level-0 message through this size-based grouping is formulated as:
@@
where $\mathcal{L}^{\text{size}}$ denotes the LLM-driven grouping operator. The resulting clusters categorize the semantic classes into groups such as SmallItems, MediumItems, and LargeItems. Details of the specific prompt used to query the LLM are provided in Appendix-I.
Next, the semantic classes within each size-based group are used as input prompts to the LLM for further clustering based on semantic relationships, with an emphasis on functional properties. This step generates the following levels structure, formulated as:
@@
where $\mathcal{L}^{\text{func}}$ denotes the LLM-driven clustering operator based on functional similarity. It is worth noting that $\{v,e\}_{1}$ does not refer strictly to a single hierarchical level. Rather, it denotes a generic stage of functional grouping that follows the initial size-based partitioning. Depending on the balance and structural design of the overall tree, the functional clustering process may be applied multiple times to generate several intermediate levels. This hierarchical refinement enables the representation to achieve maximum semantic compactness while maintaining accuracy and interpretability. The specific prompt used for this functional clustering is provided in Appendix-I.
After generating the preceding semantic substructures using LLMs across multiple levels, 3D generative models are employed to further cluster semantic classes based on their geometric characteristics. This geometry-based symbolic representation complements the language-derived hierarchy, enabling a unified structure that incorporates both semantic and geometric information. Specifically, we adopt MeshGPT [23] to capture geometric information for each semantic class. This text-to-3D framework encodes object shapes and internal topological relationships into latent, quantized embeddings, resulting in a symbolic 3D representation within the geometric shape space. Given a text prompt (i.e., the class name) as input to the pre-trained MeshGPT model, we extract the latent quantized embeddings to represent the geometric information of the class:
$\bm{z}=\mathcal{E}\{g\}\in\mathbb{N}^{F\times 2}$ (4)
where $\mathcal{E}$ denotes the pre-trained MeshGPT encoder, and $\bm{z}$ is the quantized geometric embedding extracted for class $g$. Here, $F$ represents the number of triangle faces used in the mesh representation [23]. Since the number of triangle faces $F$ varies across semantic objects, we first compute an average shape embedding $\bm{z}_{\text{avg}}\in\mathbb{R}^{1\times 2}$ for each class. These averaged embeddings are then concatenated and normalized to a standard distribution (i.e., zero mean and unit variance) to ensure consistency across classes:
@@
where $\mathcal{N}$ and $\mathcal{C}$ denote the normalization and concatenation operations, respectively. The resulting normalized embeddings $\tilde{\bm{z}}$ are subsequently clustered using the K-means++ algorithm [70], denoted by the operator $\mathcal{K}$. To address the lack of descriptive labels for the resulting clusters, we further employ an LLM-based summarization module to interpret and refine the clustering outcomes:
@@
where $\mathcal{L}^{\text{sum}}$ denotes the use of an LLM to generate semantic summaries for each geometric cluster. The resulting descriptive labels, such as “box”, “flat”, or “soft”, capture the dominant geometric attributes within each group, thereby enhancing the interpretability of the hierarchical structure. Details of the LLM prompts and summarization process are provided in Appendix-I.
Whenever LLMs are employed, we ensure that the outputs are balanced, meaningful, and descriptively coherent through carefully designed prompt inputs, as detailed in Appendix-I. To further ensure the structural correctness within each level of the hierarchical tree, we introduce a loop-based critic operation performs multiple iterations. In each iteration, an LLM is followed by a validator module that checks the quality of the generated clusters. Specifically, within each LLM-based layer, the validator assesses the clustering results in terms of completeness and semantic consistency by comparing them against the input prompt. Correctly grouped nodes are retained, while erroneously included classes are removed. Any omitted classes are compiled into a new prompt and fed into the next iteration of LLM-based clustering. This process repeats until no omitted classes remain at the current level, ensuring a complete and semantically accurate hierarchical structure. Finally, a comprehensive validation process is applied to the entire hierarchical structure, including an automated check to detect and resolve duplicates or omissions across levels, followed by a manual inspection to ensure correctness and conceptual consistency. This ensures a fully balanced and semantically coherent tree.
@@
Figure 3: Visualization of the hierarchical tree structure and semantic embedding. For tree representation, each semantic class is expressed hierarchically as $g=\{v_{l},e_{m}\}$, where edges $e_{m}$ are depicted as lines connecting tree nodes. The nodes within the same level are illustrated via the same color. For tree generation, we utilize both LLMs and 3D Generative Models to extract and group messages. For hierarchical semantic embedding, we represent semantic information using a hierarchical symbolic path. We propose two types of representations: the one-hot representation and the binary representation, which can reduce the original dimensionality by up to $O(\log N)$.
III-B Hierarchical Semantic Coding and Optimization
@@
Based on the generated hierarchical tree structure, semantic information is encoded as compact, symbolic hierarchical codes and integrated into the 3D Gaussian primitives, which are learned in an end-to-end fashion from the input image stream. To ensure accurate and structured semantic understanding, we introduce a semantic hierarchical loss, composed of both Intra-level and Inter-level components. These losses are incorporated into the overall optimization during the online SLAM process. The following sections provide a detailed description of the hierarchical semantic encoding and the proposed optimization framework.
III-B1 Hierarchical Semantic Coding
As introduced in Section III-A, we model the semantic space as a hierarchical tree structure. Within this hierarchical space, each semantic class is represented in a structured and compact form that preserves essential semantic information. The proposed hierarchical semantic coding enhances the joint optimization of the SLAM system and offers valuable scalability for semantic understanding in the complex environments. Specifically, we introduce two types of hierarchical semantic coding strategies, both of which compress the original flat semantic labels while retaining crucial semantic messages.
@@ III-B1 Hierarchical Semantic Coding
One-hot representation: As illustrated in Fig. 3, the hierarchical semantic embedding $\bm{h}$ is constructed by concatenating one-hot embeddings from all levels of the tree:
@@
where subscript denotes the $l$-th tree level, and $\mathcal{C}$ represents the concatenation operator. At each level, $\bm{h}_{l}$ is an $n$-dimensional one-hot vector (as shown in colored boxes in Fig. 3) used to represent symbolic nodes (colored circles) within that level. The dimension $n$ corresponds to the maximum number of nodes at level $l$, indexing node representations from 00 to $n-1$. The total dimensionality of the hierarchical semantic embedding is given by $N=\sum n$, i.e., the sum over all levels. Compared to the original flat representation of dimension $N_{\text{original}}$, this hierarchical structure enables an information-preserving reduction of up to $O(\log N)$ in representational complexity. The resulting compact semantic code significantly reduces memory usage and facilitates efficient joint optimization in the SLAM system.
Binary representation: To further improve compactness, we propose a binary version of the hierarchical semantic representation. The binary embedding $\bm{b}$ is constructed by concatenating the binary vectors $\bm{b}_{l}$ across all levels of the hierarchical tree:
$\bm{b}=\mathcal{C}(\bm{b}_{l})\in\mathbb{B}^{K},\quad\bm{b}_{l}\in\mathbb{B}^{k},\quad\text{where }k=\left\lceil\log_{2}n\right\rceil$ (8)
where $\bm{b}_{l}$ is the binary code converted from the one-hot embedding $\bm{h}_{l}$ at level $l$, i.e., $\bm{b}_{l}=\text{Binary}(\bm{h}_{l})$, as illustrated in the final row of Fig. 3. For example, if the one-hot vector has dimension $n=8$ (representing $8$ distinct semantic nodes indexed from 00 to $7$), its binary representation requires only $k=3$ bits for $82^{3}=$ 8. More generally, the binary code length at each level is reduced to $k=\lceil\log_{2}n\rceil$, compared to the $n$-dimensional one-hot encoding.
@@ III-B2 Loss Calculation
To effectively and efficiently optimize the proposed hierarchical semantic codings, we introduce a hierarchical loss formulation that supervises semantic embedding both within each level (intra-level) and across levels (inter-level). Specifically, the total semantic loss is defined as:
$L_{\text{{Semantic}}}=\omega_{1}L_{\text{{Intra}}}+\omega_{2}L_{\text{{Inter}}}$ (9)
where $\omega_{1}$ and $\omega_{2}$ are balancing coefficients for each loss term. Specifically, the intra-level loss $L_{\text{Intra}}$ is employed within each level’s embedding $\bm{h}^{l}$ to ensure the optimization of semantic information within each level:
$L_{\text{{Intra}}}=\sum_{l=0}^{L}L_{\text{{ce}}}(\mathcal{S}(\bm{h}^{l}),\mathcal{P}^{l})$ (10)
where $L_{\text{{ce}}}$ indicates the cross-entropy semantic loss, and $\mathcal{P}^{l}$ denotes the semantic ground truth at level $l$. $\mathcal{S}$ stands for the Softmax operation, converting embeddings into probabilities. For our proposed compact version with binary representation $\bm{b}^{l}$, we employ the binary cross-entropy loss $L_{\text{bce}}$, instead of the standard cross-entropy, for computing the binary intra-level loss:
@@
To ensure accurate semantic understanding across all levels of the hierarchical structure, we introduce the inter-level loss $L_{\text{Inter}}$, which enforces consistency between the hierarchical representation and the original flat semantic distribution. To this end, we employ a lightweight decoder to transform the holistic hierarchical embedding $\bm{h}$ into flat semantic predictions. Specifically, the decoder consists of two stacked 2D convolutional layers, with a ReLU activation applied between them. The output is then passed through a softmax function to produce a probability distribution over the flat semantic classes:
@@
The inter-level loss is defined as:
$L_{\text{Inter}}=L_{\text{ce}}\big{(}\mathcal{S}(\text{decoder}(\bm{h})),\mathcal{P}\big{)},$ (13)
where $\mathcal{S}(\cdot)$ denotes the softmax function, $\mathcal{P}$ is the ground truth flat semantic label, and $L_{\text{CE}}$ is the standard cross-entropy loss.